The angular position of a swinging door is described by θ(t)= (5.00+ 2.00t³) rad


Determine the angular position, angular speed, and angular acceleration of the door at t=3.0s.


1
Expert's answer
2022-01-16T13:12:18-0500

1) angular position:


"θ(3)= (5.00+ 2.00·3³)=59\text{ rad}."


2) angular speed:


"\omega(t)=\theta'(t)=6t^2→\omega(3)=54\text { rad/s}."


3) angular acceleration:

"\alpha(t)=\omega'(t)=12t=36\text{ rad/s}^2."